40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 7023     Entry time: Wed Jul 25 11:22:39 2012
Author: Liz 
Type: Update 
Category: Computer Scripts / Programs 
Subject: Week 6 update 

This week, I made several modifications to the Summary page scripts, made preliminary Microphone BLRMS channels and, with Rana's help, got the Weather Station working again. 

I changed the spectrogram and spectrum options in the Summary Pages so that, given the sampling frequency (which is gathered by the program), the NFFT and overlap are calculated internally.  This is an improvement over user-entered values because it saves the time of having to know the sampling frequency for each desired plot.  In addition, I set up another .sh file that can generate summary pages for any given day.  Although this will probably not be useful in the final site, it is quite helpful now because I can go back and populate the pages.  The current summary pages file is called "c1_summary_page.sh" and the one that is set up to get a specific day is called "liz_c1_summary_page.sh".  I also made a few adjustments to the .css file for the webpage so that plots completely show up (they were getting cut off on the edges before) and are easier to see.  I also figured out that the minute and second trend options weren't working because the channel names have to be modified to CHANNEL.mean, CHANNEL.min and CHANNEL.max.  So that is all in working order now, although I'm not sure if I should just use the mean trends or look at all of them (the plots could get crowded if I choose to do this).  Another modification I made to the python summary page script was adding an option to have an image on one of the pages.  This was useful because I can now put requested MEDM screens up on the site.  The image option can be accessed if, in the configuration file, you use "image-" instead of "data-" for the first word of the section header.

I also added a link to the final summary page website on the 40 meter wiki page (my summary page are currently located in the summary-test pages, but they will be moved over once they are more finalized).  I fleshed out the graphs on the summary pages as well, and have useful plots for the OSEM and OPLEV channels.  Instead of using the STS BLRMS channels, I have decided to use the GUR BLRMS channels that Masha made.  I ELOGged about my progress and asked for any advice or recommendations a few days ago (7012) and it would still be great if everyone could take a look at what I currently have up on the website and tell me what they think!  July 22 and 23 are the most finalized pages thus far, so are probably the best to look at.

https://nodus.ligo.caltech.edu:30889/40m-summary-test/archive_daily/20120723/

 

This week, I also tried to fix the problems with the Weather Station, which had not been operational since 2010.  All of the channels on the weather station monitor seemed to be producing accurate data except the rain gauge, so I went on the roof of the Machine Shop to see if anything was blatantly wrong with it.  Other than a lot of dust and spiders, it was in working condition.  I plan on going up again to clean it because, in the manual, it is recommended that the rain collector be cleaned every one to two years...  I also cleared the "daily rain" option on the monitor and set all rain-related things to zero.  Rana and I then traced the cabled from c1pem1 to the weather station monitor, and found that thy were disconnected.  In fact, the connector was broken apart and the pins were bent.  After we reconnected them, the weather station was once again operational!  In order to prevent accidental disconnection in the future, it may be wise to secure this connection with cable ties.  It went out of order again briefly on Tuesday, but I reconnected it and now it is in much sturdier shape!

 

The most recent thing that I have been doing in relation to my project has been making BLRMS channels for the MIC channels.  With Jenne's assistance, I made the channels, compiled and ran the model on c1sus, made filters, and included the channels on the PEM MEDM screen .  I have a few modifications to make and want to .  One issue that I have come across is that the sampling rate for the PEM system is 2 kHz, and the audio frequencies range all the way up to 20 kHz.  Because of this, I am only taking BLRMS data in the 1-1000 Hz range.  This may be problematic because some of these channels may only show noise (For example, 1-3 and 3-10 Hz may be completely useless).

 

The pictures below are of the main connections in the Weather Station.  This first is the one that Rana and I connected (it is now better connected and looks like a small beige box), located near the beam-splitter chamber, and the second is the c1pem1 rack.  For more information on the subject, there is a convenient wiki page: https://wiki-40m.ligo.caltech.edu/Weather_Station

Attachment 1: P7230026.JPG  3.793 MB  | Hide | Hide all
P7230026.JPG
Attachment 2: P7230031.JPG  3.946 MB  | Hide | Hide all
P7230031.JPG
ELOG V3.1.3-